A determination calls for 125cm3 of glacial acetic acid .what mass of the acid is equivalent to the required volume
a)20.8g
b)62.5g
c)116.8g
d)132.3
1
Expert's answer
2017-10-23T05:33:08-0400
Dencity of glacial acetic acid is 1.049 g/cm3
d=m/V
m=d×V=1.049 g/〖cm〗^3 ×125 〖cm〗^3=132.3 g